OpenStudy (anonymous):

The area of the sector wit hthe central angle a=120 degrees radians is equal to 24 square neters . how do i find the radius of the circle?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

degrees radians? Which is it?

OpenStudy (anonymous):

 Like this?

OpenStudy (anonymous):

I would think we'd use the following equations:\[\Large \frac{120^\circ}{360^\circ}=\frac{24m^2}{A} \]\[\Large A = \pi r^2 \]

OpenStudy (anonymous):

So I would first solve for A, then solve for r. The second equation is just the equation for the area of a circle. The first equation is just using the fact that the angle is proportional to the area of the sector.

OpenStudy (anonymous):

its 120 degrees

OpenStudy (anonymous):

so how do i solve to get the radius

OpenStudy (anonymous):

First find the area.

OpenStudy (anonymous):

120/360 pie r 2

OpenStudy (anonymous):

area us approx 603.18

OpenStudy (anonymous):

Nope! Solve for A in the first equation.

OpenStudy (anonymous):

It's going to be a round number

OpenStudy (anonymous):

i dont have r though?

OpenStudy (anonymous):

is A= pir r squared

OpenStudy (anonymous):

Use \[\Large \frac{120^\circ}{360^\circ}=\frac{24m^2}{A}\]

OpenStudy (anonymous):

Cross multiply.

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

how can you cross multuiply that?

OpenStudy (anonymous):

\[120^\circ \times A=360^\circ \times 24m^2\]

OpenStudy (anonymous):

Then divide both sides by \[120 ^\circ \]

OpenStudy (anonymous):

ohhh kk

OpenStudy (anonymous):

a=3 * 24m2

OpenStudy (anonymous):

can you walk me through the rest of the question?

OpenStudy (anonymous):

Okay, so now we can use our second equation to find the radius\[ \large A =3\times 24m^2 = 72m^2=\pi r^2 \text{ basically } 72m^2=\pi r^2 \]

OpenStudy (anonymous):

So you need to solve for \(r\).

OpenStudy (anonymous):

I'd start by dividing by \(\pi\)

OpenStudy (anonymous):

Then you'd just take the square root of both sides so that \(r^2\) becomes \(r\).

OpenStudy (anonymous):

what does the equation look like after i divide by pie

OpenStudy (anonymous):

\[\large \frac{72}{\pi}m^2 = r^2\]

OpenStudy (anonymous):

and then after i take the square root of both sides?